OpenStudy (anonymous):

If f(x) = 5x – 4 and g(x) = 6x + 3, what is f(g(x))?

OpenStudy (anonymous):

f(g(x)) = 5(6x + 3) - 4 = 30x + 15 - 4 = 30x +11

OpenStudy (anonymous):

in f(x) put 6x+3 in place of x. u get f(g(x))=5(6x+3)-4=30x+11

OpenStudy (anonymous):

Like what I did?!

OpenStudy (anonymous):

@smr_kml yeah, exactly

OpenStudy (anonymous):

Thanks @Siddharth18